A sling shot accelerates a 12g stone to a velocity of 35m/s within a distance of 5cm. To what (constsnt) force is the stone subjected during the acceleration.
1
Expert's answer
2021-08-23T12:41:24-0400

The acceleration of a stone

"a=\frac{v^2}{2d}=\frac{(35\:\rm m/s)^2}{2*0.05\:\rm m}=12250\:\rm m/s^2"

The force

"F=ma=0.012\:\rm kg*12250\:\rm m/s^2=147\: N"